# Modeling Multimodal Social Interactions: New Challenges and Baselines with Densely Aligned Representations

### v1

@@ Abstract
Understanding social interactions involving both verbal and non-verbal cues is essential to effectively interpret social situations. However, most prior works on multimodal social cues focus predominantly on single-person behaviors or rely on holistic visual representations that are not densely aligned to utterances in multi-party environments. They are limited in modeling the intricate dynamics of multi-party interactions. In this paper, we introduce three new challenging tasks to model the fine-grained dynamics between multiple people: speaking target identification, pronoun coreference resolution, and mentioned player prediction. We contribute extensive data annotations to curate these new challenges in social deduction game settings. Furthermore, we propose a novel multimodal baseline that leverages densely aligned language-visual representations by synchronizing visual features with their corresponding utterances. This facilitates concurrently capturing verbal and non-verbal cues pertinent to social reasoning. Experiments demonstrate the effectiveness of the proposed approach with densely aligned multimodal representations in modeling social interactions. We will release our benchmarks and source code to facilitate further research.
@@ 1 Introduction
Real-world social interactions involve intricate behaviors between multiple people. People communicate not only through spoken language, but also through non-verbal cues like gestures. While spoken language conveys direct meaning, inferring social meaning from it alone is sometimes vague. Non-verbal cues often play a crucial role in clarifying subtle social nuances. Therefore, comprehensively understanding social interactions involving multimodal social cues is essential to interpret social situations appropriately.
Social deduction games, where players take on roles and try to reveal the hidden roles of their opponents, provide an effective testbed for studying multimodal social interactions. These games require players to engage in communication, deception, inference, and collaboration, encompassing rich social interactions. These social interactions comprise various multimodal cues including verbal (e.g., language) and non-verbal (e.g., gesture, gaze) communications. Modeling such multimodal interactions in social deduction games can facilitate developing social artificial intelligence that can engage naturally alongside humans.
@@
There have been attempts to investigate social behaviors in multimodal aspects by considering language and visual cues together. Some works tried to learn the relationships between spoken language and visual gestures for gesture generation [1, 32, 2] and gesture-language grounding [29]. Other multimodal approaches utilized the interconnection between spoken language and visual expressions to recognize human sentiment and emotion [53, 42, 22, 39, 21, 9]. However, these works focus mainly on the behavior of a single person or rely on holistic visual representations in multi-party environments, rather than modeling the fine-grained dynamics of social interactions among multiple people.
@@
Recently, a multimodal work [28] addressed social behaviors in social deduction games where multiple players actively interact with one another. This work leveraged language and visual cues to predict persuasion strategies at the utterance level. However, it has limitations in modeling social interactions in terms of its task and methodology. Regarding the task, although persuasion strategies emerge in communication, it is largely a matter of understanding the social behavior of a single person rather than the dynamics among multiple people. In addition, it is difficult for their approach to distinguish and recognize fine-grained interactions because it utilizes holistic visual representations for the whole scene, despite multiple players being present.
@@
To address these issues, we introduce three new multimodal tasks that encompass the dynamics of player interactions, along with a corresponding baseline model. We extend the social deduction game datasets [28] with additional extensive data annotations to curate new social tasks. Our curated tasks focus on identifying referents in multi-party conversations, a critical aspect of understanding social interactions. Figure 1 shows the overall concepts of our curated tasks. The three curated tasks are as follows.
@@
Mentioned player prediction: Predicting who is mentioned by a name in a conversation.
These tasks are challenging as they require understanding the fine-grained dynamics of player interactions. For instance, when an utterance is accompanied by visual cues such as pointing gestures, it is required to figure out who is the speaker and who is being pointed at within the visual scene based on the context of utterances. This entails matching visually identified individuals with their corresponding references in spoken utterances, which enables understanding verbal and non-verbal cues comprehensively. Therefore, it is needed to align utterances with the corresponding player visuals and to utilize such densely aligned multimodal representations to tackle the tasks. Figure 2 shows an example of densely aligned representations.
To this end, we propose a novel baseline model leveraging densely aligned multimodal representations to address these new social tasks. We detect and visually track each player appearing in the video frames to distinguish individual behaviors. By establishing an initial alignment between player visual positions and their utterances, we continuously track player visuals in sync with dialogues. This alignment allows us to visually identify the speaker and the other players from given utterances. Encoding the visual gestures of a speaker and the relative positions of the other players allows us to decipher visual spatial relationships crucial for understanding non-verbal cues. Consequently, we can predict intended referents effectively by leveraging densely aligned multimodal representations from utterance language features and the corresponding visual features.
@@
We introduce new tasks in social deduction games via extensive data annotations: speaking target identification, pronoun coreference resolution, and mentioned player prediction. These tasks are challenging as they require understanding the fine-grained dynamics of interactions.
We propose a novel multimodal baseline for understanding social interactions in environments where multiple people actively interact with each other. To the best of our knowledge, this is the first work to address the dense alignment issue between language and visual social cues.
@@
| Mentioned Player Prediction | ”I’m the troublemaker and I switched [Name] with somebody” | 3,360 | – | 472 | – |
@@ 2.1 Social Behavior Analysis
Analyzing social behaviors has been widely investigated in the fields of computer vision and natural language processing. Various works have focused primarily on analyzing social behaviors from a single-modal perspective. In terms of visual cues, some works proposed gaze target estimation techniques [11, 18, 27, 48, 47] to analyze where a person is looking within a scene. There have also been studies that recognize social gaze patterns between multiple people such as identifying shared attention [20, 16, 45, 36]. Gesture recognition approaches [54, 31, 56, 30, 3] have been researched to identify specific types of human gestures such as shaking hands and thumbs-up. Regarding language cues, dialogue act recognition methods [46, 50, 40, 41, 8, 34] have been introduced to understand the communicative intent behind utterances in social dialogues. Furthermore, there have been works for sentiment analysis and emotion recognition based on dialogue language [23, 4, 44, 57, 55].
Recently, joint modeling of visual and language modalities has been studied for social behavior analysis. Some works focused on learning the relationships between spoken languages and gestures for gesture generation [1, 32, 2] and gesture-language grounding [29]. Liu et al. [32] proposed a multimodal model that integrates visual, language, and speech cues via hierarchical manners to synthesize naturalistic gestures. Additionally, the intersection of spoken utterances and visual expressions has been explored for sentiment analysis and emotion recognition [53, 42, 22, 39, 21, 9]. Hu et al. [22] proposed a unified feature space to capture the knowledge of sentiment and emotion comprehensively from multimodal cues. There also have been multimodal works for question & answering in social contexts [52, 51, 37].
However, these works mainly focus on the behaviors of a single person or rely on holistic visual features that are not densely aligned to languages in multi-party environments. They are unable to model the complex dynamics of interactions, which requires understanding the spatial relationships of multiple people in addition to their utterances. We propose a novel baseline leveraging densely aligned language-visual representations to capture the fine-grained dynamics.
@@ 2.2 Social Deduction Game Modeling
There have been works to investigate computational models for social deduction games where players actively communicate and strategize with one another. Some prior studies have focused on developing game-playing agents and analyzing optimal strategies using game theory [6, 35, 5, 43, 12]. These works aim to model the state of the game computationally but do not address understanding the dialog and behaviors of players. Chittaranjan et al. [10] modeled game outcomes from communication patterns such as player speaking and interrupting behaviors. Bakhtin et al. [14] built an agent that can play diplomacy game by utilizing language models with strategic reasoning. These approaches do not capture verbal and non-verbal multimodal aspects of modeling social behaviors. Recently, Lai et al. [28] addressed social behaviors in social deduction games using multimodal representations. They leveraged language and visual cues to predict persuasion strategies at the utterance level such as identity declaration and interrogation.
However, this multimodal work is limited in addressing multi-person dynamics due to the lack of recognizing person-level features. To address this gap, we introduce three new benchmark tasks that explicitly demand recognizing the fine-grained dynamics between individual players across both language and visual representations.
@@ 3.1 Base Datasets
Ego4D dataset. This Ego4D dataset is a subset of Ego4D Social dataset [19]. It has 40 games of One Night Ultimate Werewolf and 8 games of The Resistance: Avalon. It contains 101 separate videos with 7.3 hours. Among them, we leverage 83 videos where we can visually identify individuals for new data annotations. To guarantee the visibility of all players within the frame, this dataset adopts third-person view videos instead of first-person view videos. It also consists of videos, transcripts, player roles, voting outcomes, and persuasion strategy labels. The transcripts contain 5,815 utterances during the game.
@@ 3.2 Data Annotation
To address the fine-grained dynamics of interactions, we design three new tasks in social deduction games: speaking target identification, pronoun coreference resolution, and mentioned player prediction. Annotators reference both transcripts and videos comprehensively to make their annotations in the transcripts. To achieve reliable quality of annotations, we initially request three annotators to label subsets of data and measure their annotation agreement using Krippendorff’s alpha [26]. After we train the annotators sufficiently with larger than 0.8, we request the three annotators to label independently for the rest of the data. Note that Krippendorff’s alpha 0.8 is generally considered to have good reliability with a high level of agreement [7].
Table 1 shows the summary of our data annotation results. We achieve sufficiently high for both speaking target identification and pronoun coreference resolution. Note that we obtain the annotations for the mentioned player prediction automatically. In the training and testing process, we anonymized all names in transcripts for our tasks (e.g., [“David”, “Alice”, “Thomas”] [“Player1”, “Player2”, “Player3”]). For each task, a test set is constructed using about 20% of the annotations. We split the training and test sets at the video level rather than at the individual utterance level to ensure no overlap in terms of contextual utterances. Annotation details for each task are as follows.
Task1: Speaking target identification. This task is to predict who a speaker is talking to in a dialogue. To this end, we annotate the labels of who a speaker is speaking to at the utterance level. Since utterances are often directed to all players, we performed the annotation on only the utterances that include [“you”, “your”]. We give our annotators three labeling options: (To Player#), (To Everyone), and (To Unknown). Based on the annotations, we added “(To [MASK])” at the end of a target utterance.
Task2: Pronoun coreference resolution. This task is to predict who a pronoun refers to in a dialogue. We conduct the annotation on the third-person pronouns that are used in our dataset, which are [“he”, “she”, “him”, “her”, “his”] in utterances. We give two options to annotators: [Player#] and [Unknown]. We only target the pronouns that indicate a specific player in the game. In terms of modeling the task, we changed a pronoun into [MASK] in a target utterance.
Task3: Mentioned player prediction. This task is to predict who is referred to by their name in a dialogue. Since we know the ground truth names in utterances, it is possible to annotate these mentioned player labels automatically. We mask a mentioned player name with [MASK] in utterances and predict who is referred to in that part.
@@ 4 Proposed Approach
The introduced social tasks can be formulated as follows. Let denote utterance texts that include the -th target utterance with a [MASK] token representing an unknown player, while indicates the corresponding video frames aligned with the utterance timeline. Given and , our objective is to optimize the multimodal predictive function to effectively classify the identity of the player associated with the [MASK].
@@ 4.1 Language-Visual Alignment
A key challenge in multimodal analysis for social interactions is establishing fine-grained alignments between visual and language cues. To address this, we introduce a technique to obtain densely aligned multimodal representations from utterances and the corresponding visual cues.
We detect and track players visually in video frames overtime using AlphaPose framework [17]. Once we initially match player visuals with the player references in the utterances, we can continuously identify players in both visual and language domains. This allows us to match who is speaking in the transcript with the location of the speaker in the video, and to understand where the listeners are located relative to the speaker. By employing language-visual alignment, we can achieve densely aligned representations that integrate both verbal and non-verbal cues. It enables us to tackle our social tasks effectively, allowing for a more nuanced and holistic understanding of interactions.
@@ 4.2 Visual Interaction Modeling
To distinguish individual players on video frames , we use the human pose keypoints from AlphaPose. Specifically, we extract 17 body keypoints (, ) for each player. Figure 3 shows the procedure of encoding visual interactions. The upper path (green) of Figure 3 indicates encoding a kinesics feature of a speaker while the middle path (purple) represents encoding spatial positions of players.
First, we use the keypoints of a speaker in the upper path of Figure 3. Among 17 keypoints, we leverage [nose, l-eye, r-eye, l-shoulder, r-shoulder, l-elbow, r-elbow, l-wrist, r-wrist] which are closely related to gaze and gesture characteristics. Let , denote the image coordinates of a part at time . For example, , indicates the nose point. To represent human motion in a unified coordinate, we normalize speaker keypoints by subtracting the speaker nose point from each part point. Each point vector , is independently encoded by an MLP point encoder into part point feature ( is channel dim). These part point features are concatenated and processed by an MLP kinesics encoder to obtain a speaker kinesics feature as follows.
@@
In the meantime, the middle path (purple) of Figure 3 receives the position of each player. We consider the nose point of each player as their representative position. We normalize their nose points by subtracting the speaker’s nose point from them to get their relative positions from the speaker. We utilize their representative positions at a single time step corresponding to the start of the utterance. Let , denote the representative position of player #. Each point vector , is independently fed to an MLP point encoder to get a player point feature . We concatenate the player point features and feed them to an MLP position encoder to get . We then make aware of speaker knowledge. To this end, we obtain a speaker-label feature by passing a speaker-label one-hot vector through an FC layer. We combine with to obtain player position feature . These procedures are formulated as follows.
@@
Based on the speaker kinesics features and player position feature , we encode the visual interaction by capturing speaker kinesics motion with the context of player visual positions. and are passed through a visual interaction encoder sequentially which has the form of the transformer [49]. allows modeling dependencies between the speaker kinesics and player positions across time via self-attention. Finally, we can obtain visual interaction features that represent dynamics between the speaker and players based on the kinesics and their positions.

### v2

@@ Abstract
Understanding social interactions involving both verbal and non-verbal cues is essential for effectively interpreting social situations. However, most prior works on multimodal social cues focus predominantly on single-person behaviors or rely on holistic visual representations that are not aligned to utterances in multi-party environments. Consequently, they are limited in modeling the intricate dynamics of multi-party interactions. In this paper, we introduce three new challenging tasks to model the fine-grained dynamics between multiple people: speaking target identification, pronoun coreference resolution, and mentioned player prediction. We contribute extensive data annotations to curate these new challenges in social deduction game settings. Furthermore, we propose a novel multimodal baseline that leverages densely aligned language-visual representations by synchronizing visual features with their corresponding utterances. This facilitates concurrently capturing verbal and non-verbal cues pertinent to social reasoning. Experiments demonstrate the effectiveness of the proposed approach with densely aligned multimodal representations in modeling fine-grained social interactions. Project website: https://sangmin-git.github.io/projects/MMSI.
@@ 1 Introduction
Real-world social interactions involve intricate behaviors between multiple people. People communicate not only through verbal cues (e.g., language) but also through non-verbal cues (e.g., gesture, gaze). While spoken language conveys explicit meaning, inferring the full social context from language alone can sometimes be ambiguous. Non-verbal cues can often play a crucial role in clarifying these subtle social nuances and providing additional context. Consequently, comprehensively understanding social interactions involving multimodal social cues is essential to interpret social situations appropriately.
@@
There have been attempts to investigate social behaviors in multimodal aspects by jointly considering language and visual cues. Some works tried to learn the relationships between spoken language and visual gestures for gesture generation [1, 32, 2] and gesture-language grounding [29]. Other multimodal approaches utilized the interconnection between spoken language and visual expressions to recognize human sentiment and emotion [55, 43, 22, 39, 21, 9]. However, these works mainly focus on single-person behaviors or rely on holistic visual representations in multi-party settings, rather than modeling the fine-grained dynamics of social interactions among multiple people. Modeling such intricate multi-party dynamics requires understanding the relationships between language and visual cues at an individual level, rather than using global representations.
@@
Recently, a multimodal work [28] addressed social behaviors in social deduction games, which provide an effective testbed for studying multimodal social interactions. In these games, players take on roles and attempt to deduce the roles of their opponents by engaging in communication, deception, inference, and collaboration. These games encompass rich social interactions including verbal and non-verbal cues in multi-party settings. Lai et al. [28] leveraged language and visual cues to predict persuasion strategies at the utterance level. However, their work has limitations in modeling multi-party interactions in terms of its task and methodology. Although persuasion strategies emerge in communication, the task primarily focuses on understanding the social behaviors of a single person rather than the dynamics among people. Moreover, their approach is limited in distinguishing and recognizing fine-grained interactions because it utilizes holistic visual representations for the entire scene, despite the presence of multiple people.
@@
To address these issues, we introduce three new multimodal tasks that encompass the dynamics of multi-party interactions, along with a novel baseline model. We extend the social deduction game datasets [28] with extensive data annotations to curate new social tasks focused on identifying referents in multi-party conversations. Appropriately identifying referents is important because it significantly affects interpreting the social intent and context of conversations. Figure 1 shows the overall concepts of our curated social tasks. The three tasks are as follows.
@@
Mentioned player prediction: Predicting who is mentioned by name in a conversation.
These tasks are challenging as they require understanding the fine-grained dynamics of player interactions. For instance, when an utterance is accompanied by visual cues like pointing gestures, it is necessary to comprehensively interpret the utterance content and the speaker’s gestures for holistic reasoning. To this end, we need to figure out who the speaker is and who is being pointed at within the visual scene based on the utterance context. This involves matching utterances with the visually identified individuals. Thus, it is required to align language references with player visuals and to derive densely aligned multimodal representations from such aligned features. Figure 2 shows the concept of densely aligned language-visual representations.
To this end, we propose a novel baseline model leveraging densely aligned language-visual representations. We detect and visually track each player in the video to distinguish individual players. By initially aligning player visual positions with their language references, we continuously track player visuals in sync with the dialogue. This alignment allows the model to visually identify the speaker and the other players (i.e., listeners) for given utterances. Encoding speaker gestures and the relative positions of the listeners enables deciphering visual relationships for understanding non-verbal dynamics. We then combine this with the linguistic context of the conversation to obtain densely aligned representations. Consequently, we can effectively predict referents by leveraging densely aligned representations containing both verbal and non-verbal dynamics.
@@
We introduce new social tasks in multi-party settings via extensive data annotations: speaking target identification, pronoun coreference resolution, and mentioned player prediction. These tasks are challenging as they require understanding the fine-grained dynamics of interactions.
We propose a novel multimodal baseline model leveraging language and visual cues for understanding multi-party social interactions. To the best of our knowledge, this is the first work to address the multimodal dense alignment between language and visual social cues.
@@
| Mentioned Player Prediction | I’m the troublemaker and I switched [Name] with somebody | 3,360 | N/A | 472 | N/A |
@@ 2.1 Social Behavior Analysis
Analyzing social behaviors has been widely investigated in the fields of computer vision and natural language processing. Various works have focused primarily on analyzing social behaviors from a single-modal perspective. In terms of visual cues, some works proposed gaze target estimation techniques [11, 18, 27, 49, 48] to analyze where a person is looking within a scene. There have also been studies that recognize social gaze patterns between multiple people such as identifying shared attention [20, 16, 46, 36]. Gesture recognition approaches [56, 31, 58, 30, 3] have been researched to identify specific types of human gestures such as shaking hands and thumbs-up. Regarding language cues, dialogue act recognition methods [47, 51, 40, 41, 8, 34] have been introduced to understand the communicative intent behind utterances in social dialogues. Furthermore, there have been works on sentiment analysis and emotion recognition based on dialogue language [23, 4, 45, 59, 57].
Recently, joint modeling of visual and language modalities has been studied for social behavior analysis. Some works focused on learning the relationships between spoken language and gestures for gesture generation [1, 32, 2] and gesture-language grounding [29]. Liu et al. [32] proposed a multimodal model that integrates visual, language, and speech cues via hierarchical manners to synthesize naturalistic gestures. Additionally, the intersection of spoken utterances and visual expressions has been explored for sentiment analysis and emotion recognition [55, 43, 22, 39, 21, 9]. Hu et al. [22] proposed a unified feature space to capture the knowledge of sentiment and emotion comprehensively from multimodal cues. There have also been multimodal works for question & answering in social contexts [54, 53, 37].
However, these works mainly focus on the behaviors of a single person or rely on holistic visual features that are not densely aligned to language in multi-party environments. They are unable to model the complex dynamics of interactions, which requires understanding the spatial relationships of multiple people in addition to their utterances. We propose a novel approach leveraging densely aligned language-visual representations to capture the fine-grained dynamics.
@@ 2.2 Social Deduction Game Modeling
There have been works investigating computational models for social deduction games where players actively communicate and strategize with one another. Some prior studies have focused on developing game-playing agents and analyzing optimal strategies using game theory [6, 35, 5, 44, 12]. These works aim to model the state of the game computationally but do not address understanding the dialogue and behaviors of players. Chittaranjan et al. [10] modeled game outcomes from communication patterns such as player speaking and interrupting behaviors. Bakhtin et al. [14] built an agent that can play diplomacy games by utilizing language models with strategic reasoning. These approaches do not capture verbal and non-verbal multimodal aspects of modeling social behaviors. Recently, Lai et al. [28] addressed social behaviors in social deduction games using multimodal representations. They leveraged language and visual cues to predict persuasion strategies at the utterance level such as identity declaration and interrogation.
However, this multimodal work is limited in addressing multi-party dynamics due to the lack of person-level feature recognition. To address this gap, we introduce new tasks in social deduction games that explicitly demand recognizing person-level features. We also propose the corresponding baseline model that captures the multi-party dynamics across both language and visual representations.
@@ 3.1 Base Datasets
Ego4D dataset. This Ego4D dataset is a subset of Ego4D Social dataset [19]. It has 40 games of One Night Ultimate Werewolf and 8 games of The Resistance: Avalon. It contains 101 separate videos with 7.3 hours. Among them, we leverage 83 videos where we can visually identify individuals for new data annotations. To guarantee the visibility of all players within the frame, this dataset adopts third-person view videos instead of first-person view videos. It also consists of videos, transcripts, player roles, voting outcomes, and persuasion strategy labels. The transcripts contain 5,815 utterances during the games.
@@ 3.2 Data Annotation
To address the fine-grained dynamics of interactions, we design three new tasks in social deduction games: speaking target identification, pronoun coreference resolution, and mentioned player prediction. Annotators reference both transcripts and videos comprehensively to make their annotations in the transcripts. To achieve reliable annotation quality, we initially request three annotators to label subsets of data and measure their annotation agreement using Krippendorff’s alpha [26]. After we train the annotators sufficiently with larger than 0.8, we request the three annotators to label independently for the rest of the data. Note that Krippendorff’s alpha 0.8 is generally considered to indicate good reliability with a high level of agreement [7].
Table 1 shows the summary of our data annotation results. We achieve sufficiently high for both speaking target identification and pronoun coreference resolution. Note that we could obtain the annotations for the mentioned player prediction automatically. In the training and testing process, we anonymize all names in transcripts (e.g., [“David”, “Alice”, “Thomas”] [“Player1”, “Player2”, “Player3”]). For each task, a test set is constructed using about 20% of the annotations. We split the training and test sets at the video level rather than at the individual utterance level to ensure no overlap in terms of contextual utterances. Annotation details for each task are as follows.
Task1: Speaking target identification. This task aims to predict who a speaker is talking to in a dialogue. To this end, we annotate the labels of who a speaker is speaking to at the utterance level. Since utterances are often directed to all players, we perform the annotation only on the utterances that include [“you”, “your”]. We give our annotators three labeling options: (To Player#), (To Everyone), and (To Unknown). Based on the annotations, we add “(To [MASK])” at the end of a target utterance.
Task2: Pronoun coreference resolution. This task aims to predict who a pronoun refers to in a dialogue. We conduct the annotation on the third-person pronouns that are used in our dataset, which are [“he”, “she”, “him”, “her”, “his”] in utterances. We give two options to annotators: [Player#] and [Unknown]. We only target the pronouns that indicate a specific player in the game. In terms of modeling the task, we change a pronoun to [MASK] in a target utterance.
Task3: Mentioned player prediction. This task aims to predict who is referred to by their name in a dialogue. Since we know the ground truth names in utterances, it is possible to annotate these mentioned player labels automatically. We modify a mentioned player name with [MASK] in utterances and predict who is referred to in that part.
@@ 4 Proposed Approach
The introduced social tasks can be formulated as follows. Let denote utterance texts that include the -th target utterance containing a [MASK] token representing an unknown player, while indicates the corresponding video frames aligned with the utterance timeline. Given and , our objective is to optimize the multimodal predictive function to effectively classify the identity of the player associated with the [MASK].
@@ 4.1 Language-Visual Alignment
To achieve densely aligned representations containing both verbal and non-verbal dynamics properly, it is necessary to first establish fine-grained alignments between visual and language cues. To this end, we distinguish individual visuals and match language references with them.
We detect and track players visually in video frames over time using AlphaPose framework [17]. Once we initially match player visuals with the player references in the utterances (i.e., assigning each tracking ID to Player#), we can continuously identify players in both visual and language domains. It enables the model to identify the speaker in the video based on the corresponding utterance and to determine the spatial arrangement of the listeners. Building upon this alignment, we can achieve densely aligned language-visual representations from verbal and non-verbal cues in a comprehensive manner. It enables us to tackle our social tasks effectively, allowing for a more nuanced and holistic understanding of multi-party interactions.
@@ 4.2 Visual Interaction Modeling
To distinguish individual players in video frames , we use the human pose keypoints from AlphaPose. Specifically, we extract 17 body keypoints (, ) for each player. Figure 3 shows the procedure of encoding visual interactions. The upper path (green) of Figure 3 indicates encoding a kinesics feature of a speaker, while the middle path (purple) represents encoding spatial positions of all players.
First, we use the keypoints of a speaker in the upper path of Figure 3. Among the 17 part keypoints, we leverage [nose, l-eye, r-eye, l-shoulder, r-shoulder, l-elbow, r-elbow, l-wrist, r-wrist] closely related to gaze and gesture characteristics. Let , denote the image coordinates of a part at time . For example, , indicates the nose point. To represent human motion in a unified coordinate, we normalize speaker keypoints by subtracting the speaker nose point from each part point. Each point vector , is encoded by an MLP point encoder into a part point feature ( is channel dim). These part point features are concatenated and processed by an MLP kinesics encoder to obtain a speaker kinesics feature as follows.
@@
In the meantime, the middle path (purple) of Figure 3 receives the position of each player. We consider the nose point of each player as their representative position. We normalize their nose points by subtracting the speaker’s nose point from them to get their relative positions from the speaker. We utilize their representative positions at a single time step corresponding to the start of the utterance. Let , denote the representative position of Player#. Each point vector , is independently fed to an MLP point encoder to get a player point feature . We concatenate the player point features and feed them to an MLP position encoder to get . We then make aware of speaker knowledge. To this end, we make a speaker-label feature by passing a speaker-label one-hot vector through an FC layer. We combine with to obtain a player position feature . These procedures are formulated as follows.
@@
Based on the speaker kinesics features and player position feature , we encode the visual interaction by capturing speaker kinesics motion with the context of player visual positions. and are passed through a visual interaction encoder sequentially, which has the form of the transformer [50]. allows modeling dependencies between the speaker kinesics and player positions across time via self-attention. Finally, we can obtain visual interaction features that represent dynamics between the speaker and players based on the speaker kinesics and listener positions.